determine the slope of the line through the given points. if the slope is undefined, so state.
(5,3) and (3,5)
My question is how do i do this? I know how to graph points but i dont understand it when they are the same numbers and i dont understand what the slope of the line means, or what undefined means.
thanks,
jennifer lewis

Answer by jainenderkapoor(61)

Slope of the line when two points are given can be found using the formula
Slope is given by Rise / Run
m = (y2 - y1)/ (x2 - x1) where the two given points are taken as (x1, y1) and (x2, y2)
here points are (5, 3) and (3, 5)
so slope m = (5 - 3)/(3 - 5) = 2/-2 = -1
The slope is actually the inclination of the line with the positive direction of x axis.
By undefined slope we mean that the line is perpendicular to the x axis
I hope this is clear to you.
In case you need Online Tutoring you can contact me at jainenderkapoor@gmail.com
Thanks